OpenStudy (anonymous):

What is the equation of the line that is perpendicular to -x + y = 7 and passes through (-1, -1)?

OpenStudy (anonymous):

A. -x + y = 2 B. -x – y = 2 C. x – y = 2 D. -x – y = 0 E. x + y = 0

OpenStudy (haseeb96):

do u know what is the slope in this equation -x +y =7 ? if u know then tell me

OpenStudy (anonymous):

I don't how do I find it?

OpenStudy (haseeb96):

first convert the above equation in slope intercept form (y=mx +b )

OpenStudy (anonymous):

y = -x + 7?

OpenStudy (haseeb96):

good y=mx+b since, m is slope and b is the y intercept whereas y=-x +7 here m=-1 okay your slope is -1 now apply this Perpendicular formula m x m'=-1 -1 x m'=-1 m'=1

OpenStudy (haseeb96):

have u got it ? we apply the perpendicular formula of slope because we require to find out the equation of line which is perpendicular to that equation

OpenStudy (anonymous):

So it's D?

OpenStudy (haseeb96):

no use this formula y-y' =m' (x-x') since m'=1 (x',y')=(-1.-1) and put it in the above formula and get the answer

OpenStudy (anonymous):

so y - -1 = 1 (x - -1)

OpenStudy (haseeb96):

yes y-(-1)=1(x-(-1)) y+1=1(x+1) y+1=x+1 x-y+1-1=0 x-y=0 this is your equation

OpenStudy (anonymous):

x - y = 0 is my equation?

OpenStudy (haseeb96):

yes plz again check your choices and point which is given to us in your question from your note book

OpenStudy (anonymous):

x + y = 0

OpenStudy (haseeb96):

yes this is correct answer do u know where u did mistake u wrote y=-x +7 it is in correct y=x+7

OpenStudy (haseeb96):

when we take out the slope from it it would be m=1 when m x m'=-1 1 x m' = -1 m'=-1 here m'=-1 (x',y')=(-1,-1) apply on these things on this formula and get the correct equation y-y'=m'(x-x') y-(-1)= -1(x-(-1)) y+1=-1(x+1) y+1=-x +1 y+x +1-1=0 x+y =0 x+y=0 this is your correct answer

OpenStudy (haseeb96):

@CaseyCarns look it here

OpenStudy (anonymous):

Thank you so much!

OpenStudy (haseeb96):

if u have any question u can ask me now?

OpenStudy (anonymous):

No I understand!

OpenStudy (haseeb96):

okay welcome here